OpenStudy (anonymous):

Taina, Luther, and Della are tapping a balloon to each other in the air, trying to keep it from touching the ground. The distance from Taina to Luther is 22 in. The distance from Luther to Della is 40 in. The distance from Della to Taina is 34 in. Find the measures of the three angles in the triangle.

OpenStudy (anonymous):



OpenStudy (anonymous):

Use \[a ^{2}=b ^{2}+c ^{2}-2bc \cos(A)\] Solving for A you get \[A = \cos ^{-1}( \frac{ b ^{2}+c ^{2}-a^{2} }{ 2bc })\] So the angle for Luther = \[\cos ^{-1}( \frac{ 40 ^{2}+22 ^{2}-34^{2} }{ 2(40)(22) })\]

OpenStudy (anonymous):

Go back and review the Law of cosines if youre still confused.